### v1

# Ore extensions of multiplier Hopf coquasigroups

Abstract In this paper, Ore extensions of multiplier Hopf coquasigroups are studied. Necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given. Then the isomorphism between two Ore extensions is discussed.
Keywords multiplier Hopf coquasigroup, Ore extension.
Mathematics Subject Classification 16T05 $\cdot$ 16T99
1 Introduction
The Ore extension is a well-known method of ring extensions to construct noncommutative rings and algebras. In Hopf algebra theory, Ore extension were also served as a key tool to create noncommutative and non-cocommutative Hopf algebras, such as Panov [9], Beattie [1, 2] and Nenciu [7, 8].
Over time, Hopf algebras has been evolved into broader frameworks. In the 1990s, Van Daele [10] pioneered the theory of multiplier Hopf algebras, breaking the unit constraint of traditional Hopf algebras and offering new tools for quantum groups and noncommutative algebra. Subsequently, its theory was gradually expanded to duality theory, representation theory, and quantum applications. Meanwhile, the theory of Hopf coquasigroups advanced through the groundbreaking work of Klim and Majid [6]. They introduced weak coassociativity to resolve dual construction challenges in non-coassociative structures. Their work also revealed deep connections with algebraic topology.
Subsequently, some algebraists investigated the feasibility of Ore extensions on these structures (Figure 1). Zhao and Lu [14] initiated research on Ore extensions of multiplier Hopf algebras, while Jiao and Meng [3] extended this to Hopf coquasigroups, proposing Hopf coquasigroup-Ore extensions.
[…]
Figure 1: Diagram of the Ore Extension Structure Relationships
Based on these advances, a natural question arises: Could Ore extensions also hold for multiplier Hopf coquasigroups? This is the main motivation.
This paper is organized as follows. In section 2, we introduce some basic notions: Ore extension, multiplier algebras and multiplier Hopf coquasigroups, which will be used in the following sections.
In section 3, we present the concept of Ore extensions in the context of multiplier Hopf coquasigroups. Moreover, we present the main theorem of this paper (Theorem 3.4), which gives a necessary and sufficient condition for Ore extensions of regular multiplier Hopf coquasigroups to be regular multiplier Hopf coquasigroups. This theorem gives a positive answer to the above question.
In section 4, we give the notion of isomorphism for Ore extensions, and show sufficient conditions for two such extensions to be isomorphic.
2 Preliminaries
Throughout this paper, all linear spaces we considered are over a fixed field $k$ (e.g., the complex number field $\mathds{C}$).
2.1 Multiplier Hopf coquasigroups
Let $A$ be an (associative) algebra over $k$, with or without identity, but the product is nondegenerate. This means that, whenever $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, we must have that $a=0$.
Recall from [10, 11] that a left multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=\rho(a)b$ for all $a,b\in A$. Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=a\rho(b)$ for all $a,b\in A$. A multiplier of $A$ is a pair $(\rho_{1},\rho_{2})$ of a left and a right multiplier such that $\rho_{2}(a)b=a\rho_{1}(b)$ for all $a,b\in A$. For any algebra $A$, there are a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, which are respectively denoted by $L(A)$, $R(A)$ and $M(A)$. In particular, $M(A)$ is characterized as the largest algebra with identity containing $A$ as an essential two-sided ideal. Additionally, we still have that, whenever $a\in M(A)$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, again $a=0$. Furthermore, we consider the tensor algebra $A\otimes A$. It is still nondegenerate and we have its multiplier algebra $M(A\otimes A)$. There are natural imbeddings
$A\otimes A\subseteq M(A)\otimes M(A)\subseteq M(A\otimes A).$
In general, when $A$ has no identity, these two inclusions are strict. If $A$ already has an identity $1_{A}$, the product is obviously nondegenerate and $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
Let $A$ and $B$ be nondegenerate algebras, if homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (i.e., $f(A)B=B$ and $Bf(A)=B$), then it has a unique extension to a homomorphism $M(A)\longrightarrow M(B)$, we also denote it $f$.
A regular multiplier Hopf coquasigroup introduced in [13] is a nondegenerate associative algebra $A$ equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$(coproduct), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode) such that
  1. (1)
    $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$ for any $a,b\in A$.
  2. (2)
    The counit satisfies $(\varepsilon\otimes id)T_{1}(a\otimes b)=ab=(id\otimes\varepsilon)T_{2}(a\otimes b)$.
  3. (3)
    $S$ is antimultiplicative and anticomultiplicative such that for any $a,b\in A$
    $(m\otimes\iota)(S\otimes\Delta)\Delta(a)=1_{M(A)}\otimes a=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(a),$ (2.1)
    $(\iota\otimes m)(\Delta\otimes S)\Delta(a)=a\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(a).$ (2.2)
Multiplier Hopf coquasigroup $(A,\Delta)$ is called regular, if the antipode $S$ is bijective. $T_{1}$ and $T_{2}$ in the above are usually called Galois maps.
Multiplier Hopf coquasigroups introduced in[13] generalized Hopf coquasigroups to a non-unital case. Unlike Hopf coquasigroups in [5, 6], the underlying algebra of multiplier Hopf coquasigroup $A$ need not have an identity in general, and the comultiplication $\Delta$ does not map $A$ into $A\otimes A$ but rather into its multiplier algebra $M(A\otimes A)$. Multiplier Hopf coquasigroups weaken the coassociativity of multiplier Hopf algebra, and unify the notions of multiplier Hopf algebras and Hopf coquasigroups.
We use the adapted Sweedler notation introduced in [12] for the comultiplication: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
2.2 Ore extention
Ore extension of an algebra
First, let us recall the definition of Ore extension of an algebra from [4]. Let $A$ be an algebra, and $\tau{}$ be an algebra endomorphism of $A$. A linear endomorphism $\delta{}$ of $A$ is called an $\tau$-derivation $\delta$ of $A$ if for all $a,b\in A$,
$\delta(ab)=\delta(a)b+\tau(a)\delta(b)$
The above condition implies $\delta(1_{A})=0$.
The Ore extension $A[y;\tau,\delta]$ of an algebra $A$ is an algebra generated by the algebra $A$ with the relation
$ya=\tau(a)y+\delta(a)$
for all $a\in A$.
Ore extension of Hopf algebra
Let $A$ and $A[y;\tau,\delta]$ be Hopf algebras over $k$. The Hopf algebra $A[y;\tau,\delta]$ is called the Hopf-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a Hopf subalgebra in $A[y;\tau,\delta]$.
Following Theorem 1.3 in [9], the necessary and sufficient conditions for the Hopf algebra $A[y;\tau,\delta]$ to be a Hopf-Ore extension are given as follows.
(1) there is a character $\chi:A\rightarrow k$ such that
$\tau(a)=\chi(a_{(1)})a_{(2)}\qquad\text{for all $a\in A$};$
(2) the following relation holds:
$\chi(a_{(1)})a_{(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)});$
(3) the $\tau$-derivation $\delta$ satisfies the relation
$\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
Ore extension of Hopf coquasigroups
Let $A$ and $A[y;\tau,\delta]$ be Hopf coquasigroups, recall from [3] that the Hopf coquasigroup $A[y;\tau,\delta]$ is a Hopf coquasigroups-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
From Theorem 3.3 in [3], for a Hopf coquasigroup $A$, the Hopf coquasigroup $H=A[y,\tau,\delta]$ becomes an Hopf coquasigroup-Ore extension if and only if the following conditions hold:
(1) There exists a character $\chi:A\rightarrow k$, such that for every $a\in A$, satisfying
$\alpha(a)=\chi(a_{(1)})a_{(2)}.$
(2) The subsequent relations hold:
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\text{Ad}_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}.$
(3) For the $\alpha$-derivation $\delta$, the relation
$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$
holds for all $a\in A$.
Ore extension of Multiplier Hopf algebras
Let $(A,\Delta)$ be a regular multiplier Hopf algebra and $A[y;\tau,\delta]$ be an Ore extension of the algebra $A$.
If $\tau$ is a nondegenerate algebra endomorphism, then $\tau$-derivation $\delta$ uniquely extends to the multiplier algebra $M(A)$ (i.e., $\overline{\delta}:M(A)\rightarrow M(A)$). If $\tau$ is surjective, following Proposition 2.3 of Section 2 in [14], the product of $A[y,\tau,\delta]$ is nondegenerate. There are some natural extensions: the comultiplication $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, the antipode $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, the counit $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$.
If restricting these extensions to the subspace $A[y;\tau,\delta]$, i.e., $\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, $\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k$, we get the comultiplication $\Delta^{\prime}$, counit $\varepsilon^{\prime}$, and the antipode $S^{\prime}$ on the Ore extension $(A[y;\tau,\delta],\Delta^{\prime})$ of $A$.
$(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra $A$, if $(A[y;\tau,\delta],\Delta^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1+r\otimes y$ for some $r\in M(A)$.
For a regular multiplier Hopf algebra $A$ with Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective. $A[y;\tau,\delta]$ constitutes a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
(1) There exists a character $\chi:A\rightarrow k$ such that, for any $a\in A$, the following equality holds:
$\tau(a)=(\chi\otimes id)\Delta(a)=r\big{(}(id\otimes\chi)\Delta(a)\big{)}r^{-1}.$
(2) For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes id)\Delta(a)+(r\otimes 1)\big{(}(id\otimes\delta)\Delta(a)\big{)}$
is satisfied.
As introduced in Section 2.1 of this paper, in the following, we also denote these unique extensions $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ as $\Delta$, $S$, $\varepsilon$ without confusion.
3 Ore extensions of multiplier Hopf coquasigroups
Let $(A,\Delta)$ be a regular multiplier Hopf coquasigroup introduced in [13]. The aim of this section is to show a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to be a regular multiplier Hopf coquasigroup.
First, let us introduce the definition of Ore extension of regular multiplier Hopf coquasigroups, which genelized the Ore extension of Hopf coquasigroups and multiplier Hopf algebras introduced in [3, 14].
Definition 3.1. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called the Ore extension of multiplier Hopf coquasigroups (or MHC-Ore extension in short) if $\Delta(y)=y\otimes r_{1}+r_{2}\otimes y$ for some $r_{1},r_{2}\in M(A)$.
Remark 3.2. (1) As in [14], $y\otimes 1+r\otimes y$ in above definition belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$.
(2) If the coproduct $\Delta$ of multiplier Hopf coquasigroup $A$ is coassociative, then $A$ is a multiplier Hopf algebra and the Ore extension is actually the one introduced in [14].
(3) If the underlying algebra of multiplier Hopf coquasigroup $A$ has an identity, then $A$ is a Hopf coquasigroup and the Ore extension is actually the one introduced in [3].
By the definition of regular multiplier Hopf coquasigroup, for $y\in A[y;\tau,\delta]$, we have
$(m\otimes\iota)(S\otimes\Delta)\Delta(y)=1_{M(A)}\otimes y=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(y),$
$(\iota\otimes m)(\Delta\otimes S)\Delta(y)=y\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(y),$
that is
$S(y)r_{1(1)}\otimes r_{1(2)}+S(r_{2})y\otimes r_{1}+S(r_{2})r_{2}\otimes y=1_{M(A)}\otimes y$
$=yS(r_{1(1)})\otimes r_{1(2)}+r_{2}S(y)\otimes r_{1}+r_{2}S(r_{2})\otimes y,$
$y\otimes S(r_{1})r_{1}+r_{2}\otimes S(y)r_{1}+r_{2(1)}\otimes S(r_{2(2)})y=y\otimes 1_{M(A)}$
$=y\otimes r_{1}S(r_{1})+r_{2}\otimes yS(r_{1})+r_{2(1)}\otimes r_{2(2)}S(y).$
Then
$\hskip 40.00006pt{r_{1}}S({r_{1}})=S({r_{1}}){r_{1}}=1,\qquad{r_{2}}S({r_{2}})=S({r_{2}}){r_{2}}=1,$
$S(y){r_{1(1)}}\otimes{r_{1(2)}}+S({r_{2}})y\otimes{r_{1}}=yS({r_{1(1)}})\otimes{r_{1(2)}}+{r_{2}}S(y)\otimes{r_{1}}=0,$
${r_{2}}\otimes S(y){r_{1}}+{r_{2(1)}}\otimes S({r_{2(2)}})y={r_{2}}\otimes yS({r_{1}})+{r_{2(1)}}\otimes{r_{2(2)}}S(y)=0.$
Let $r_{1}$ and $r_{2}$ be group-like elements in $M(A)$, then we have
$\Delta(r_{i})=r_{i}\otimes r_{i},\quad\varepsilon(r_{i})=1,\quad S(r_{i})=r_{i}^{-1},\quad i=1,2,$
$\big{(}S(y)r_{1}+S(r_{2})y\big{)}\otimes r_{1}=\big{(}yS(r_{1})+r_{2}S(y)\big{)}\otimes r_{1}=0,$
$r_{2}\otimes\big{(}S(y)r_{1}+S(r_{2})y\big{)}=r_{2}\otimes\big{(}yS(r_{1})+r_{2}S(y)\big{)}=0.$
By substituting the generator $y$ with $y^{\prime}=yr_{1}^{-1}$, it follows that the coproduct map satisfies $\Delta(y^{\prime})=y^{\prime}\otimes 1+r\otimes y^{\prime}$, where $r=r_{2}r_{1}^{-1}$ is a group-like element in the multiplier algebra $M(A)$. In the subsequent discussion, we consistently assume that the element $y$ in the Ore extension of a regular multiplier Hopf coquasigroup, the following relation holds:
$\Delta\left(y\right)=y\otimes 1+r\otimes y$ (3.1)
for some group-like element $r\in M(A)$. As usual, $A{d_{r}}\left(a\right)=raS\left(r\right)=ra{r^{-1}}$.
Now we give the following properties for the Ore extension of a regular multiplier Hopf coquasigroup.
Proposition 3.3. When $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$, then
$\varepsilon(y)=0,\quad S(y)=-r^{-1}y,\quad S^{-1}(y)=-\tau(r)^{-1}y,$
$\hskip 60.00009pt\delta(r)=0,\quad\delta(r^{-1})=0.$
proof. By applying the main property of counit, where $a=a_{(1)}\varepsilon(a_{(2)})$, it follows that $y=\varepsilon(y)+\varepsilon(r)y=\varepsilon(y)+y$, this implies $\varepsilon(y)=0$.
By applying the main property of antipode, where $\varepsilon(a)=a_{(1)}S(a_{(2)})$, it follows that $0=\varepsilon(y)=y+rS(y)$, this implies $S(y)=-r^{-1}y$. Since
$y=S^{-1}S(y)=S^{-1}(-r^{-1}y)=-S^{-1}(r^{-1}y)=-S^{-1}(y)S^{-1}(r^{-1})=-S^{-1}(y)r,$
there is an element $a\in A$ such that $S^{-1}(y)=ay$. And applying the preceding formula we get
$y=-ayr=-a\tau(r)y-a\delta(r),$
this implies that $a=-\tau(r)^{-1}$ and $\delta(r)=0$, so we get that $S^{-1}(y)=-\tau(r)^{-1}y$. Since $\delta$ is a $\tau-derivation$, we have
$0=\delta(1)=\delta(rr^{-1})=\delta(r)r^{-1}+\tau(r)\delta(r^{-1}),$
[…]
Next, we will present the main result of this section, which gives the necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup.
Theoerm 3.4. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. Assume that $\tau$ is a surjective endomorphism on $A$. Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
(1) There exists a character $\chi$: $A\rightarrow k$ satisfying
$\tau(a)=\chi(a_{(1)})a_{(2)};$ (3.2)
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)};$ (3.3)
for any $a\in A$.
(2) The $\tau-derivation$ $\delta$ satisfies the relation
$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$ (3.4)
for any $a\in A$.
Proof. The proof will be conducted in three steps.
step1. Comultiplication.
Similar to the contruction in [14], assume that the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau.\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$. Then the algebra homomorphism $\Delta$ preserves the relation $ya=\tau(a)y+\delta(a)$ for any $a\in A$, i.e.,
$\Delta(y)\Delta(a)=\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}.$
For any $a,b\in A$, we have that
$\Delta(y)\Delta(a)(1\otimes b)$
$=(y\otimes 1+r\otimes y)(a_{(1)}\otimes{a_{(2)}}b)$
$=y{a_{(1)}}\otimes{a_{(2)}}b+r{a_{(1)}}\otimes y{a_{(2)}}b$
$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}}b)y+r{a_{(1)}}\otimes\delta({a_{(2)}}b)$
$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
$\quad+r{a_{(1)}}\otimes\tau({a_{(2)}})\delta(b)$
$=\big{(}\tau({a_{(1)}})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
$\quad+\big{(}r{a_{(1)}}{r^{-1}}\otimes\delta({a_{(2)}})\big{)}(r\otimes b)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\delta(b)\big{)}+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
and
$\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(y\otimes b+r\otimes yb)+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(y\otimes b)+\Delta\big{(}\tau(a)\big{)}(r\otimes yb)+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}$
$\quad+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
So we have that $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following formulas hold:
$\Delta\big{(}\tau(a)\big{)}(1\otimes b)$ $=\tau({a_{(1)}})\otimes{a_{(2)}}b,$
$\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}$ $={r{a_{(1)}}\otimes\tau({a_{(2)}}b)},$
$\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$ $=r{a_{(1)}}\otimes\delta({a_{(2)}}b)+\delta({a_{(1)}})\otimes{a_{(2)}}b.$
By the definition of $\delta\otimes\iota$ and $\iota\otimes\delta$ in [14, Lemma 2.5] and the extensions of $\tau\otimes\iota$, the above conditions change to the following:
$\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
$\Delta\big{(}\tau(a)\big{)}=A{d_{r}}(a_{(1)})\otimes\tau(a_{(2)}),$ (3.6)
$\Delta\big{(}\delta(a)\big{)}=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)}).$
[…]
$\chi(a)b=m(\iota\otimes S)\big{(}(b\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
for any $a,b\in A$. Because $\tau$ is nondegenerate, for $b$ there are $a_{i}$ , $b_{i}$ such that $\sum{{a_{i}}}\tau\left({{b_{i}}}\right)=b$. Hence, the above definition is changed to the following:
$\chi(a)b$ $=m(id\otimes S)\Big{(}\sum\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\Big{(}\sum(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)\Big{(}(\tau\otimes id)(b_{i}\otimes 1)\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}\bigg{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)(\tau\otimes id)\Big{(}(b_{i}\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}\bigg{)}.$
Hence $\chi(a)b\in A$ for any $a,b\in A$. In addition, for any $c\in A$,
$\chi(a)(bc)$ $=m(id\otimes S)\big{(}(bc\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=m(id\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=b\big{(}m(id\otimes S)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=b\big{(}\chi(a)c\big{)}.$
[…]
$\chi(a)b=m(id\otimes S)\Big{(}(b\otimes 1)(\tau\otimes id)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
and since $b$ is arbitrary, this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}.$ Therefore, $\chi(A)\subseteq k$. We can regard $\chi$ as a mapping $\chi:A\rightarrow k$. Considering that $\tau$ is an algebra endomorphism and
$\chi(ab)=\varepsilon\big{(}\tau(ab)\big{)}=\varepsilon\big{(}\tau(a)\big{)}\varepsilon\big{(}\tau(b)\big{)}=\chi(a)\chi(b),$
$\chi(a+b)=\varepsilon\big{(}\tau(a+b)\big{)}=\varepsilon\big{(}\tau(a)\big{)}+\varepsilon\big{(}\tau(b)\big{)}=\chi(a)+\chi(b),$
[…]
$(id\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
[…]
$S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}=1\otimes\tau(a)=\tau(a_{(1)})S(a_{(2)(1)})\otimes a_{(2)(2)}.$ (3.7)
By applying the formula $\chi\left(a\right)=\varepsilon\left(\tau\left(a\right)\right)$, $\tau$ can be derived from $\chi$,
$\chi(a_{(1)})a_{(2)}$ $=\varepsilon\big{(}\tau(a_{(1)})\big{)}a_{(2)}=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon(a_{(2)(1)})a_{(2)(2)}$
$=\varepsilon\big{(}\tau(a_{(1)})\big{)}\varepsilon\big{(}S(a_{(2)(1)})\big{)}a_{(2)(2)}$
$=\varepsilon\big{(}\tau(a_{(1)})\big{)}S(a_{(2)(1)})a_{(2)(2)}$
$=\varepsilon(1)\tau(a)=\tau(a).$
[…]
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)},$
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=A{d_{r}}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}.$
[…]
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}=\tau(a_{(1)})\otimes a_{(2)},$
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}=Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
[…]
Therefore, we can obtain that the comultiplication of $A$ (i.e., the comultiplication on $M(A)$ ) can be extended to an algebra homomorphism $\Delta:M(A)[y;\tau,\delta]\rightarrow M(A)[y;\tau,\delta]\otimes M(A)[y;\tau,\delta]$, and $\Delta$ does not necessarily have to be coassociative.
Step2. Counit.
Assume that $A[y;\tau,\delta]$ already has the comultiplication defined as in the above step. In what follows, we will prove the existence of a counit denoted by $\overline{\varepsilon}$ extending $\varepsilon$ from $A$(in fact $M(A)$) to the Ore extension $M(A)[y;\tau,\delta]$. Define $\overline{\varepsilon}(y)=0$. We need to verify the equation
$\overline{\varepsilon}(y)\varepsilon(x)=\varepsilon\big{(}\tau(x)\big{)}\overline{\varepsilon}(y)+\varepsilon\big{(}\delta(x)\big{)}$
for any $x\in M(A)$, i.e., $\varepsilon\big{(}\delta(x)\big{)}=0$. In fact, we only have to prove the above formula is true for any $a\in A$ in the sense that
$0=\varepsilon\big{(}\delta(xa)\big{)}=\varepsilon\big{(}\delta(x)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}.$
[…]
$\delta(a)$ $=m(id\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
$=m(id\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
$=\delta(a_{(1)})\varepsilon(a_{(2)})+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}$
$=\delta(a)+ra_{(1)}\otimes\varepsilon\big{(}\delta(a_{(2)})\big{)}.$
Therefore, $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0$
This proves $\varepsilon\big{(}\delta(a)\big{)}=0$ , for any $a\in A$. Furthermore, since
$0=\varepsilon\big{(}\delta(xa)\big{)}$ $=\varepsilon\big{(}\delta(x)a+\tau(x)\delta(a)\big{)}$
$=\varepsilon\big{(}\delta(a)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}$
$=\varepsilon\big{(}\delta(x)\big{)}\Big{(}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\Big{)},$
it follows that for any $x\in M(A)$, $\varepsilon\big{(}\delta(x)\big{)}=0$ holds true. Consequently, $\varepsilon$ can be extended to $M(A)[y;\tau,\delta]$.
For all $\sum a_{i}y^{i},\sum b_{j}y^{j}\in A[y;\tau,\delta]$, it is easy to prove
$(\overline{\varepsilon}\otimes\iota)\big{(}\Delta(\sum a_{i}y^{i})(1\otimes\sum b_{j}y^{j})\big{)}$ $=(\sum a_{i}y^{i})(\sum b_{j}y^{j}),$
$(\iota\otimes\overline{\varepsilon})\big{(}(\sum b_{j}y^{j}\otimes 1)\Delta(\sum a_{i}y^{i})\big{)}$ $=(\sum b_{j}y^{j})(\sum a_{i}y^{i}).$
Step3. Antipode.
[…]
$(m\otimes\iota)(S\otimes\Delta)\Delta(ya)=1\otimes ya=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(ya),$ (3.8)
$(\iota\otimes m)(\Delta\otimes S)\Delta(ya)=ya\otimes 1=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(ya),$ (3.9)
and
$S(a)S(y)=S(y)S\big{(}\alpha(a)\big{)}+S\big{(}\delta(a)\big{)}$ (3.10)
[…]
$1\otimes\delta(a)$ $=(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})\big{)}\otimes\Delta(a_{(2)})+S(ra_{(1)})\otimes\Delta\big{(}\delta(a_{(2)})\big{)}\Big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\delta({a_{(1)}})\big{)}\otimes a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)})\otimes\delta(a_{(2)(1)})\otimes a_{(2)(2)}$
$\quad+S(ra_{(1)})\otimes ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
$=S\big{(}\delta(a_{(1)})\big{)}a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)})\delta(a_{(2)(1)})\otimes a_{(2)(2)}$
$\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)}).$ (3.11)
Applying the formula mentioned above, we can obtain
$\big{(}S(b)\otimes 1\big{)}(m\otimes\iota)(S\otimes\Delta)\Delta(ya)$
$=(m\otimes\iota)(S\otimes\Delta)\big{(}\Delta(ya)(b\otimes 1)\big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\Delta\big{(}\tau(a)y+\delta(a)\big{)}(b\otimes 1)\Big{)}$
$=(m\otimes\iota)(S\otimes\Delta)\bigg{(}\Big{(}\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}\Big{)}(b\otimes 1)\bigg{)}$
$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau({a_{(1)}})yb\otimes{a_{(2)}}+\tau({a_{(1)}})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
$=(m\otimes\iota)\Big{(}S\big{(}\tau(a_{(1)})yb\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}\otimes{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}\otimes a_{(2)(1)}r\otimes{a_{(2)(2)}}y\Big{)}$
$\quad+(m\otimes\iota)(S\otimes\Delta)\big{(}\delta({a_{(1)}})b\otimes{a_{(2)}}+r{a_{(1)}}b\otimes\delta({a_{(2)}})\big{)}$
$=S\big{(}\tau({a_{(1)}})yb\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
$\quad+S(r{a_{(1)}}b)\otimes\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)\otimes r{a_{(2)(1)}}\otimes\delta({a_{(2)(2)}})\Big{)}$
$=S(b)S(y)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+S\big{(}\delta(a)b\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
$\quad+S(r{a_{(1)}}b)\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)r{a_{(2)(1)}}\otimes\delta({a_{(2)(1)}})$
[…]
$\quad+(S(b)\otimes 1)\Big{(}S\big{(}\delta(a_{(1)})\big{)}{a_{(2)(1)}}\otimes a_{(2)(2)}+S(r{a_{(1)})\delta(a_{(2)(1)}})\otimes{a_{(2)(2)}}$
$\quad+S(ra_{(1)})ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
=(3.11)${=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$ )
$=S(b)\otimes\tau(a)y+S(b)\otimes\delta(a)$
$=S(b)\otimes\big{(}\tau(a)y+\delta(a)\big{)}$
$=S(b)\otimes ya$
$=\big{(}S(b)\otimes 1\big{)}(1\otimes ya).$
[…]
$-S(a)r^{-1}y$ $=-r^{-1}yS\big{(}\tau(a)\big{)}+S\big{(}\delta(a)\big{)}$
$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)},$
for any $a\in M(A).$
[…]
$S(a)r^{-1}=r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)},$ (3.12)
$\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}=rS\big{(}\delta(a)\big{)}.$ (3.13)
for any $a\in M(A).$
[…]
In fact,
$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $=\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
$=\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
$=\chi(a_{(1)})S(a_{(2)(1)})Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
$=Ad_{r}\big{(}S(a)\big{)}$
[…]
$S\big{(}\tau(a)\big{)}=S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)}.$ (3.14)
[…]
Applying $m(\iota\otimes S)$ to the equation
$(b\otimes 1)\Delta\big{(}\delta(a)\big{)}=b\delta(a_{(1)})\otimes a_{(2)}+bra_{(1)}\otimes\delta(a_{(2)}),$
it follows that
$m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\delta(a_{(1)})S(a_{(2)})+bra_{(1)}S(\delta\big{(}{a_{(2)}})\big{)}.$
By applying the properties of the antipode, we have
$m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\delta(a)\big{)}\Big{)}=b\varepsilon\big{(}\delta(a)\big{)}=0,$
i.e,
$bra_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-b\delta(a_{(1)})S(a_{(2)})\qquad\text{for any $a,b\in A$}.$
And considering the nondegeneracy of the product of $A$, this means that the following equation holds for any $a\in A$
$a_{(1)}S\big{(}\delta(a_{(2)})\big{)}=-r^{-1}\delta(a_{(1)})S(a_{(2)}).$ (3.15)
Hence, we have
$rS\big{(}\delta(a)\big{)}$ $=rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
=(3.15)${=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
$=-Ad_{r}S({{a_{(1)}}})\delta(a_{(2)(1)})S(a_{(2)(2)}),$ (3.16)
for any $a\in A$.
In addition, from the action of $\delta$ on the element $\varepsilon(a)=a_{1}S(a_{(2)})$, it follows that
$0=\delta(a_{(1)})S(a_{(2)})+\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)},$
i.e.,
$\tau(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=-\delta(a_{(1)})S(a_{(2)}).$ (3.17)
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $=\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
$=\chi(a_{(1)})\tau\big{(}{S(a_{(2)(1)}})\big{)}\tau(a_{(2)(2)(1)})\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
=(3.17)${=}-\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})\big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$ )
$=-\tau\Big{(}S\big{(}\chi(a_{(1)})a_{(2)(1)}\big{)}\Big{)}\delta(a_{(2)(2)(1)})S(a_{(2)(2)(2)})$
=(3.3)${=}-\tau\Big{(}S\big{(}\chi(a_{(1)(1)})a_{(1)(2)}\big{)}\Big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
$=-\chi(a_{(1)(1)})\tau\big{(}S(a_{(1)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-\chi(a_{(1)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\chi\big{(}S(a_{(1)(2)(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-\chi(a_{(1)(1)})S(a_{(1)(2)(1)})Ad_{r}\big{(}S(a_{(1)(2)(2)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-Ad_{r}\big{(}S(a_{(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$
=(3.16)${=}rS\big{(}\delta(a)\big{)}$ )
for any $a,b\in A$.
[…]
4 Isomorphism of Ore extensions
From Theorem 3.4, it follows that a MHC-Ore extension is determined by $\chi$, $r$, and $\delta$. Thus, $A(\chi,r,\delta)$ is used to denote the MHC-Ore extension of $A[y;\tau,\delta]$, where $\chi$ is a character of $A$, $r$ is a group-like element in $M(A)$ and $\delta$ is a $\langle\chi,r\rangle$-derivation. The corresponding Ore extension on the multiplier coquasigroup is denoted by $M(A)(\chi,r,\delta)$.
Now we can define an isomorphism of MHC-Ore extensions.
Definition 4.1. Two MHC-Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ of regular multiplier Hopf coquasigroups $A$ and $A^{\prime}$ are said to be isomorphic if there is an isomorphism of multiplier Hopf coquasigroups $\phi$: $A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ such that $\phi(A)=A^{\prime}$.
Similarly to [14, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
Theorem 4.2. For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with coproducts: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$. Then $A(\chi,r,\delta)$ is isomorphic to $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ if there is an isomorphism $\phi:A\rightarrow A^{\prime}$ such that
$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},$ (4.1)
where $d^{\prime}\in A^{\prime}$ such that $\Delta(d^{\prime})=d^{\prime}\otimes 1+r^{\prime}\otimes d^{\prime}$.
Proof. Let $\phi(y)=y^{\prime}+d^{\prime}$. Then $\phi$ can be extended from $\phi:A\rightarrow A^{\prime}$ to $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$. We have
$\phi(ya)$ $=\phi\big{(}\tau(a)y+\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}\phi\big{(}y\big{)}+\phi\big{(}\delta(a)\big{)}$
$=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
$\phi(y)\phi(a)$ $=(y^{\prime}+d^{\prime})\phi(a)=\tau^{\prime}\big{(}\phi(a)\big{)}y^{\prime}+\delta^{\prime}\big{(}\phi(a)\big{)}+d^{\prime}\phi(a)$
$=\tau^{\prime}\big{(}\phi(a)\big{)}y^{\prime}+\phi\big{(}\delta(a)\big{)}+\phi\big{(}\tau(a)\big{)}d^{\prime}-d^{\prime}\phi(a)+d^{\prime}\phi(a)$
$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
for any $a\in A$. Thus we get
$\phi(ya)=\phi(y)\phi(a).$
Since
$\Delta\big{(}\phi(y)\big{)}$ $=\Delta(y^{\prime}+d^{\prime})=\Delta(y^{\prime})+\Delta(d^{\prime})$
$=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}+d^{\prime}\otimes 1+r^{\prime}\otimes d^{\prime}$
$=(y^{\prime}+d^{\prime})\otimes 1+r^{\prime}\otimes(y^{\prime}+d^{\prime}),$
$(\phi\otimes\phi)\Delta(y)$ $=(\phi\otimes\phi)(y\otimes 1+r\otimes y)$
$=(y^{\prime}+d^{\prime})\otimes 1+r^{\prime}\otimes(y^{\prime}+d^{\prime}),$
we can obtain
$\Delta\big{(}\phi(y)\big{)}=(\phi\otimes\phi)\Delta(y).$
Considering the properties of $\varepsilon$, we obtain
$d^{\prime}=(\varepsilon\otimes id)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
this indicates $\varepsilon(d^{\prime})=0$. And by Proposition 3.3, we know that $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. So we have
$\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y^{\prime}+d^{\prime})=\varepsilon(y^{\prime})+\varepsilon(d^{\prime})=0,$
this means that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
Considering the properties of $S$, we obtain
$0=\varepsilon(d^{\prime})=m(S\otimes id)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
this indicates $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Thus we have
$\phi\big{(}S(y)\big{)}=\phi(-{r^{-1}}y)=-\phi({r^{-1}})\phi(y)=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
$S\big{(}\phi(y)\big{)}=S(y^{\prime}+d^{\prime})=S(y^{\prime})+S(d^{\prime})=-{r^{-1}}y^{\prime}-r{{}^{\prime-1}}d^{\prime}=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
it follows that $\phi\big{(}S(y)\big{)}=S\big{(}\phi(y)\big{)}$.
[…]
Acknowledgements
The work was partially supported by the China Postdoctoral Science Foundation (No. 2019M651764) and National Natural Science Foundation of China (No. 11601231).